A student measures a single chocolate cookie and finds that it has an average diameter of 5.24±0.02cm. The cookie is near to a perfect cylinder and has a width of 5.5±0.2mm.


Calculate

i. The volume of the cookie 


i. The density, assuming if the student measures the cookie on a top pan balance to be 10g±1.5 %. Give your answer in g/cm and also in kg/m³


1
Expert's answer
2021-08-29T16:55:18-0400

1)


"V=0.25\pi d^2h\\V=0.25\pi (5.24)^2(0.55)=12\ cm^3"

2)


"\rho=\frac{10}{12}=0.83\frac{g}{cm^3}\\\rho=830\frac{kg}{m^3}"